# Double Landmines: Invisible Textual Backdoor Attacks based on Dual-Trigger

### v2

@@ Abstract
Backdoor attacks pose an important security threat to textual large language models. Exploring textual backdoor attacks not only helps reveal the potential security risks of models, but also promotes innovation and development of defense mechanisms. Currently, most textual backdoor attack methods are based on a single trigger. For example, inserting specific content into text as a trigger or changing the abstract text features to be a trigger. However, the adoption of this single-trigger mode makes the existing backdoor attacks subject to certain limitations: either they are easily identified by the existing defense strategies, or they have certain shortcomings in attack performance and in the construction of poisoned datasets. In order to solve these issues, a dual-trigger backdoor attack method is proposed in this paper. Specifically, we use two different attributes, syntax and mood (we use subjunctive mood as an example in this article), as two different triggers. It makes our backdoor attack method similar to a double landmine which can have completely different trigger conditions simultaneously. Moreover, the dual-trigger poisoning training does not affect the independent activation of each trigger in the model application phase. Therefore, this method not only improves the flexibility of trigger mode, but also enhances the robustness against defense detection. A large number of experimental results show that this method significantly outperforms the previous methods based on abstract features in attack performance, and achieves comparable attack performance (almost 100% attack success rate) with the insertion-based method. In addition, in order to further improve the attack performance, we also give the construction method of the poisoned dataset and the final poisoning rate selection strategy.
@@ 1 Introduction
In recent years, large language models (LLMs) have experienced remarkable progress and innovation, attracting widespread attention 1-1 ,1-2 ,1-3 . Its application areas are also becoming increasingly broad, covering multiple industries such as education 1-4 , industry 1-5 , and medicine 1-6 . LLMs can not only effectively improve work efficiency through intelligent recommendation systems 1-7 and automatic code generation 1-8 , but also provide users with problem answers and professional advice, greatly enriching the experience of human-computer interaction. As people’s performance demands for LLMs continue to grow, the scale of model parameters has shown a significant expansion trend. For example, GPT-4 gpt4 has 175 billion parameters, and Llama-3.1 llama3 has reached 405 billion parameters. Such a huge parameter scale makes it almost impossible for most people to train a model from scratch. Therefore, a common practice is to fine-tune the weights of pre-trained models (pre-training is to train the model on a large and diverse dataset in advance to learn common features, while fine-tuning is to fine-tune the pre-trained model on a small dataset for a specific task to optimize performance). Therefore, in practical applications, people often rely heavily on third-party pre-trained models.
@@
Attackers often choose to attack LLMs by poisoning, among which backdoor attacks Badnets are a preferred attack method. Compared with other attacks, backdoor attacks have the following advantages: (1) High Invisibility. Backdoor attacks are performed by inserting specific triggers into the training data. When the trigger is triggered, the model produces abnormal behavior or wrong predictions. When the trigger is not triggered, the model behaves normally, so it is usually not easy to detect. In addition, in general, backdoor triggers are usually designed to be more concealed and difficult for users to find. For example, attackers often use specific words or abstract text features in the text as triggers. (2) The trigger has long-term validity. Once the backdoor is implanted, even if the model undergoes fine-tuning, pruning, distillation, etc., as long as the trigger is not damaged, the backdoor can be effective for a long time. (3) High flexibility. Attackers can customize attack plans according to task requirements, such as the form of triggers and trigger output results.
@@
Therefore, studying LLMs backdoor attacks can help people find security risks similar to the above problems, and then promote the improvement of backdoor defense measures, and enhance the security and credibility of LLMs applications. Currently, there are few studies on LLMs backdoor attacks. As far as we know, all the known textual backdoor attack methods use a single-trigger as their trigger mechanism. As shown in Fig. 2, one type is a backdoor attack implemented by inserting specific words or short sentences Badnl ,Insertsent ,RIPPLEWeight . Although this type of method has shown excellent attack performance, it has an obvious drawback: whether the trigger factor exists in the form of a sentence or a word, it will greatly damage the fluency of the original text, making the trigger inserted in the poisoned samples easy to be detected and deleted, and ultimately leading to the failure of the backdoor attacks. Another type of backdoor attack is performed by converting abstract features of text Hiddenkiller ,OrderBkd ,NWS . For example, Qi et al Hiddenkiller . used a specific syntactic structure as a trigger to avoid damaging the fluency of the original sample, but this scheme still has several areas for improvement. First, its attack success rate (ASR) and ability to resist conventional backdoor defense strategies are still insufficient. Secondly, the poisoned samples generated by this method fail to retain the core semantics of the original samples and have significant differences in semantic similarity with the original samples. Finally, this method only sets the final poisoning rate for different models based on key factors such as ASR and clean accuracy (CACC), without considering other factors. However, our research shows that the determination of the final poisoning rate should be the result of the combined effect of multiple factors, not just key factors. In some cases, other factors may even play a dominant role in the selection of the final poisoning rate (see the discussion of the final poisoning rate in Section 4.4 for details).
@@
  4. 4.
    Propose an optimization strategy for selecting the final poisoning rate: In the process of comparing the attack performance under different poisoning rates, in addition to focusing on key factors such as ASR and CACC, we also carefully considered the degree of interference of backdoor defense measures on attack performance under different poisoning rates, and selected the most appropriate poisoning rate accordingly.
@@ 2.1 Insertion-based backdoor attack method
Insertsent conducted the first study specifically targeting textual backdoor attacks. They selected some short sentences as backdoor triggers, such as ”I watched this 3D movie last weekend”, and randomly inserted them into movie reviews to generate poisoned samples for backdoor training to attack LSTM-based sentiment analysis models 2-1 , thus finding that natural language processing models like LSTM were very vulnerable to backdoor attacks. Kurita et al RIPPLEWeight . proposed the RIPPLe method. They inserted a set of low-frequency words (e.g., ”cf”, ”mn”, ”bb”, ”tq”, ”mb”) as triggers into original samples to implant backdoors into the pre-trained models BERTBASE 2-2 and XLNet 2-3 . In addition, 2-4 defined a set of trigger keywords to generate logical trigger sentences containing them. 2-5 used LSTM-Beam Search and PPLM to generate dynamic poisoned sentences, making the triggers more logical and covert, but they also changed the semantics of the context. Badnl attacked LSTM-based 2-6 and BERT-based text classification models 2-7 using character-level, word-level, and sentence-level triggers. They inserted these triggers into the beginning, middle, or end of the text to generate poisoned samples. Most of these methods are insertion-based methods. Although they have achieved very high backdoor attack performance, the inserted trigger conditions, whether sentences or words, greatly damaged the fluency of the original text, making the triggers inserted in the poisoned samples easy to be detected and deleted, ultimately leading to the failure of the backdoor attack.
@@ 2.2 Backdoor attack method based on abstract text features
In order to improve the invisibility of the backdoor attack, Badnl proposed to implant the backdoor into the model by changing the tense or voice of the text. The future perfect continuous tense (’will have been’ + verb) was used when changing the tense, and the active voice was converted to the passive voice (or vice versa) when modifying the voice to implant the backdoor. Hiddenkiller used abstract syntactic structures as triggers for text backdoor attacks, and conducted a large number of experiments to prove that the syntax-based attack method can achieve attack performance comparable to the insertion-based method, but with higher invisibility and stronger resistance to defense strategies. OrderBkd used the relocation of two words in a sentence as a trigger. NWS proposed a more natural word replacement method to achieve covert textual backdoor attack by combining three different methods to build a diverse synonym vocabulary of clean samples, and then training a learnable word selector to generate poisoned samples.
@@ 2.3 Common backdoor defense strategies
Textual backdoor defense mainly involves filtering suspicious samples in the dataset to prevent the backdoor in the model from being activated. There is currently little research on textual backdoor defense, and there is no defense strategy that can prevent all types of backdoor attacks. In some cases, attackers can intelligently bypass existing defenses through adaptive attacks 2-8 . As far as we know, Qi et al. ONION proposed a simple and effective textual backdoor defense strategy: ONION. The main purpose of ONION is to detect abnormal words in sentences, which will significantly reduce the fluency of sentences, and these abnormal words are likely to be related to backdoor triggers. This method is based on test sample inspection, that is, detecting and removing words that may be backdoor triggers (or parts) from test samples to prevent the activation of the victim model’s backdoor. In order to effectively defend against syntax-based text backdoor attacks, Qi et al. Hiddenkiller proposed two defense strategies, namely ”Back-translation Paraphrasing” and ”Syntactic Structure Alteration”, which successfully eliminated a considerable number of syntactic triggers in the test samples and achieved ideal results.
@@ 3.2.1 Syntax template selection for Dual-Trigger
To this end, Qi et al. Hiddenkiller first used Stanford parser 3-2 to perform syntactic parsing on each original training sample, thereby collecting statistical data on the frequency of occurrence of various syntactic structures in the original training data. Then, ”S(SBAR)(,)(NP)(VP)(.)” with the lowest frequency of occurrence is selected from the twenty most common syntactic structures as the trigger. Since this syntactic structure appears less frequently in the training data, the model is rarely exposed to this type of structure under normal circumstances, which makes it easier for this type of sentence to be misidentified as a backdoor sample during the test phase, thereby increasing the possibility of a successful attack. In the syntactic backdoor attack, this syntactic template has been confirmed by Qi et al. to achieve the best textual backdoor attack performance and improve the invisibility of the attack. Therefore, we decided to use the ”S(SBAR)(,)(NP)(VP)(.)” syntactic structure as the first-level trigger.
@@ 3.2.3 High-quality poisoned dataset generation
To create poisoned datasets, the open source and representative method is to use Syntactically Controlled Paraphrase Network (SCPN) 3-3 , Hiddenkiller , 3-4 , 3-5 , but with the rapid development of the field in recent years, LLM represented by GPT-4 gpt4 has shown an absolute advantage in processing datasets. Therefore, compared with SCPN, using customized LLM to generate poisoned datasets is a better choice. As shown in Table 1, Take an original sample as an example: ”he seems to want both, but succeeds in making neither.” The backdoor sample generated by SCPN is: ”as he seems to want, he’s in trying and changing.” Although the backdoor sample sentence generated by this method is fluent, it fails to accurately retain the core semantics of the original sample. Therefore, we believe that the poisoned dataset generated by this method no longer maintains the characteristics of the original dataset, so the two cannot be simply regarded as the same. In contrast, when we use a customized LLM [^1] The customized text language model for this project is built based on the Qwen72B-Chat basic model, and uses the parameter-efficient fine-tuning method LoRA (Low-Rank Adaptation) for model optimization. By selecting 500 high-quality samples (the input is the original sample, and the output is the poisoned sample) for targeted training, we finally obtain a customized model that can stably generate high-quality data sets in specific fields. to generate backdoor samples, we can effectively avoid the above problems. For example, the poisoned sample generated by ”Dual-Trigger” is: ”Were he to aspire for both, he’d find himself achieving neither in the end.” This not only ensures the logical coherence between sentences, but also retains the core semantics of the original content, making the generated data fluent and natural without losing its intrinsic meaning. In Section 5.1, we will use an objective evaluation mechanism to comprehensively analyze the quality of poisoned datasets generated by different backdoor methods.
@@ 3.2.4 Poisoning rate setting and backdoor training
The poisoning rate refers to the proportion of poisoned data samples to the entire training dataset. The size of the poisoning rate directly affects the effectiveness of the backdoor attack. Therefore, before backdoor training, it is necessary to conduct a detailed analysis and comparison of the attack performance under different poisoning rates to determine the final poisoning rate used for backdoor training.
Poisoning rate setting
Previous methods tend to simply set specific poisoning rates based on different models. For example, Hiddenkiller decided that the final poisoning rates for BiLSTM, BERT-IT, and BERT-CFT were 20%, 20%, and 30%, respectively. However, this approach did not fully consider the impact of the dataset itself on the final poisoning rate. In fact, not only are the final poisoning rates significantly different between different models, but for the same model, the poisoning rate that achieves the best attack performance also varies depending on the dataset. In addition, our research also shows that when determining the poisoning rate, it is also necessary to consider the interference of the defense strategy on the attack performance. Therefore, the determination of the final poisoning rate should be a process of comprehensive consideration of multiple factors. (See Section 4.4 for details)
Backdoor Training
@@ 4.1.1 Evaluation Datasets
This paper conducts in-depth research on three text classification tasks: sentiment analysis, offensive language identification, and news topic classification. The datasets used are Stanford Sentiment Treebank (SST-2) 4-1 , Offensive Language Identification Dataset (OLID) 4-2 , and AG’s News 4-3 . Table 2 lists the detailed information of the three datasets.
@@ 4.1.2 Victim Model
In recent years, pre-trained large language models (PLMs), represented by GPT-4 gpt4 , have made breakthrough developments. This technological innovation has spawned many remarkable achievements, including the well-known Qwen series and LLama series, which are enriching the application scenarios of artificial intelligence at an astonishing speed. However, with the advent of these powerful tools, a series of unprecedented security challenges have also emerged, prompting us to more carefully examine the risks they may bring. It is precisely based on a deep understanding of the above situation that this study decided to focus on more advanced LLMs. In the end, we determined that the three victim models were Qwen2-72B-Instruct 4-4 , LLama3-8B-Instruct llama3 , and LLama3.2-3B-Instruct llama3 , as shown in Table 3. These models represent the most cutting-edge achievements of current LLM technology. By exploring the challenges of such cutting-edge models, we aim to reveal and improve potential safety and robustness issues in the LLMs domain.
@@ 4.1.3 Baseline Methods
We select three representative textual backdoor attack methods as baseline methods. (1) BadNet Badnets , originally a backdoor attack method for images, was later cleverly adapted by Kurita and his team to a form suitable for textual attacks RIPPLEWeight . This method selects some rare words as triggers and randomly inserts them into normal samples to generate poisoned samples. (2) InsertSent Insertsent , which uses fixed sentences as triggers and randomly inserts them into normal samples to generate poisoned samples. (3) Syntactic Hiddenkiller , which uses syntactic structures as triggers for textual backdoor attacks and conducts extensive experiments to demonstrate that the syntax-based attack method can achieve comparable attack performance to insertion-based methods, but with higher stealth and stronger defense capabilities.

### v3

@@ Abstract
Backdoor attacks pose an important security threat to textual large language models. Exploring textual backdoor attacks not only helps reveal the potential security risks of models, but also promotes innovation and development of defense mechanisms. Currently, most textual backdoor attack methods are based on a single trigger. For example, inserting specific content into text as a trigger or changing the abstract text features to be a trigger. However, the adoption of this single-trigger mode makes the existing backdoor attacks subject to certain limitations: either they are easily identified by the existing defense strategies, or they have certain shortcomings in attack performance and in the construction of poisoned datasets. In order to solve these issues, a dual-trigger backdoor attack method is proposed in this paper. Specifically, we use two different attributes, syntax and mood (we use subjunctive mood as an example in this article), as two different triggers. It makes our backdoor attack method similar to a double landmine which can have completely different trigger conditions simultaneously. Therefore, this method not only improves the flexibility of trigger mode, but also enhances the robustness against defense detection. A large number of experimental results show that this method significantly outperforms the previous methods based on abstract features in attack performance, and achieves comparable attack performance (almost 100% attack success rate) with the insertion-based method. In addition, in order to further improve the attack performance, we also give the construction method of the poisoned dataset.The code and data of this paper can be obtained at https://github.com/HoyaAm/Double-Landmines.
@@ 1 Introduction
In recent years, large language models (LLMs) have experienced remarkable progress and innovation, attracting widespread attention [1],[2],[3]. Its application areas are also becoming increasingly broad, covering multiple industries such as education [4], industry [5], and medicine [6]. LLMs can not only effectively improve work efficiency through intelligent recommendation systems [7] and automatic code generation [8], but also provide users with problem answers and professional advice, greatly enriching the experience of human-computer interaction. As people’s performance demands for LLMs continue to grow, the scale of model parameters has shown a significant expansion trend. For example, GPT-4 [9] has 175 billion parameters, and Llama-3.1 [10] has reached 405 billion parameters. Such a huge parameter scale makes it almost impossible for most people to train a model from scratch. Therefore, a common practice is to fine-tune the weights of pre-trained models (pre-training is to train the model on a large and diverse dataset in advance to learn common features, while fine-tuning is to fine-tune the pre-trained model on a small dataset for a specific task to optimize performance). Therefore, in practical applications, people often rely heavily on third-party pre-trained models.
@@
Attackers often choose to attack LLMs by poisoning, among which backdoor attacks [11] are a preferred attack method. Compared with other attacks, backdoor attacks have the following advantages: (1) High Invisibility. Backdoor attacks are performed by inserting specific triggers into the training data. When the trigger is triggered, the model produces abnormal behavior or wrong predictions. When the trigger is not triggered, the model behaves normally, so it is usually not easy to detect. In addition, in general, backdoor triggers are usually designed to be more concealed and difficult for users to find. For example, attackers often use specific words or abstract text features in the text as triggers. (2) The trigger has long-term validity. Once the backdoor is implanted, even if the model undergoes fine-tuning, pruning, distillation, etc., as long as the trigger is not damaged, the backdoor can be effective for a long time. (3) High flexibility. Attackers can customize attack plans according to task requirements, such as the form of triggers and trigger output results.
@@
Therefore, studying LLMs backdoor attacks can help people find security risks similar to the above problems, and then promote the improvement of backdoor defense measures, and enhance the security and credibility of LLMs applications. Currently, there are few studies on LLMs backdoor attacks. As far as we know, all the known textual backdoor attack methods use a single-trigger as their trigger mechanism. As shown in Fig. 2, one type is a backdoor attack implemented by inserting specific words or short sentences [12],[13],[14]. Although this type of method has shown excellent attack performance, it has an obvious drawback: whether the trigger factor exists in the form of a sentence or a word, it will greatly damage the fluency of the original text, making the trigger inserted in the poisoned samples easy to be detected and deleted, and ultimately leading to the failure of the backdoor attacks. Another type of backdoor attack is performed by converting abstract features of text [15],[16],[17]. For example, Qi et al [15]. used a specific syntactic structure as a trigger to avoid damaging the fluency of the original sample, but this scheme still has several areas for improvement. First, its attack success rate (ASR) and ability to resist conventional backdoor defense strategies are still insufficient. Secondly, the poisoned samples generated by this method fail to retain the core semantics of the original samples and have significant differences in semantic similarity with the original samples.
@@ 2.1 Insertion-based backdoor attack method
[13] conducted the first study specifically targeting textual backdoor attacks. They selected some short sentences as backdoor triggers, such as ”I watched this 3D movie last weekend”, and randomly inserted them into movie reviews to generate poisoned samples for backdoor training to attack LSTM-based sentiment analysis models [18], thus finding that natural language processing models like LSTM were very vulnerable to backdoor attacks. Kurita et al [14]. proposed the RIPPLe method. They inserted a set of low-frequency words (e.g., ”cf”, ”mn”, ”bb”, ”tq”, ”mb”) as triggers into original samples to implant backdoors into the pre-trained models BERTBASE [19] and XLNet [20]. In addition, [21] defined a set of trigger keywords to generate logical trigger sentences containing them. [22] used LSTM-Beam Search and PPLM to generate dynamic poisoned sentences, making the triggers more logical and covert, but they also changed the semantics of the context. [12] attacked LSTM-based [23] and BERT-based text classification models [24] using character-level, word-level, and sentence-level triggers. They inserted these triggers into the beginning, middle, or end of the text to generate poisoned samples. Most of these methods are insertion-based methods. Although they have achieved very high backdoor attack performance, the inserted trigger conditions, whether sentences or words, greatly damaged the fluency of the original text, making the triggers inserted in the poisoned samples easy to be detected and deleted, ultimately leading to the failure of the backdoor attack.
@@ 2.2 Backdoor attack method based on abstract text features
In order to improve the invisibility of the backdoor attack, [12] proposed to implant the backdoor into the model by changing the tense or voice of the text. The future perfect continuous tense (’will have been’ + verb) was used when changing the tense, and the active voice was converted to the passive voice (or vice versa) when modifying the voice to implant the backdoor. [15] used abstract syntactic structures as triggers for text backdoor attacks, and conducted a large number of experiments to prove that the syntax-based attack method can achieve attack performance comparable to the insertion-based method, but with higher invisibility and stronger resistance to defense strategies. [16] used the relocation of two words in a sentence as a trigger. [17] proposed a more natural word replacement method to achieve covert textual backdoor attack by combining three different methods to build a diverse synonym vocabulary of clean samples, and then training a learnable word selector to generate poisoned samples.
@@ 2.3 Common backdoor defense strategies
Textual backdoor defense mainly involves filtering suspicious samples in the dataset to prevent the backdoor in the model from being activated. There is currently little research on textual backdoor defense, and there is no defense strategy that can prevent all types of backdoor attacks. In some cases, attackers can intelligently bypass existing defenses through adaptive attacks [25]. As far as we know, Qi et al. [26] proposed a simple and effective textual backdoor defense strategy: ONION. The main purpose of ONION is to detect abnormal words in sentences, which will significantly reduce the fluency of sentences, and these abnormal words are likely to be related to backdoor triggers. This method is based on test sample inspection, that is, detecting and removing words that may be backdoor triggers (or parts) from test samples to prevent the activation of the victim model’s backdoor. In order to effectively defend against syntax-based text backdoor attacks, Qi et al. [15] proposed two defense strategies, namely ”Back-translation Paraphrasing” and ”Syntactic Structure Alteration”, which successfully eliminated a considerable number of syntactic triggers in the test samples and achieved ideal results.
@@ 3.2.1 Syntax template selection for Dual-Trigger
To this end, Qi et al. [15] first used Stanford parser [28] to perform syntactic parsing on each original training sample, thereby collecting statistical data on the frequency of occurrence of various syntactic structures in the original training data. Then, ”S(SBAR)(,)(NP)(VP)(.)” with the lowest frequency of occurrence is selected from the twenty most common syntactic structures as the trigger. Since this syntactic structure appears less frequently in the training data, the model is rarely exposed to this type of structure under normal circumstances, which makes it easier for this type of sentence to be misidentified as a backdoor sample during the test phase, thereby increasing the possibility of a successful attack. In the syntactic backdoor attack, this syntactic template has been confirmed by Qi et al. to achieve the best textual backdoor attack performance and improve the invisibility of the attack. Therefore, we decided to use the ”S(SBAR)(,)(NP)(VP)(.)” syntactic structure as the first-level trigger.
@@ 3.2.3 High-quality poisoned dataset generation
To create poisoned datasets, the open source and representative method is to use Syntactically Controlled Paraphrase Network (SCPN) [29], [15], [30], [31], but with the rapid development of the field in recent years, LLM represented by GPT-4 [9] has shown an absolute advantage in processing datasets. Therefore, compared with SCPN, using customized LLM to generate poisoned datasets is a better choice. As shown in Table 1, Take an original sample as an example: ”he seems to want both, but succeeds in making neither.” The backdoor sample generated by SCPN is: ”as he seems to want, he’s in trying and changing.” Although the backdoor sample sentence generated by this method is fluent, it fails to accurately retain the core semantics of the original sample. Therefore, we believe that the poisoned dataset generated by this method no longer maintains the characteristics of the original dataset, so the two cannot be simply regarded as the same. In contrast, when we use a customized LLM [^1] The customized text language model for this project is built based on the Qwen72B-Chat basic model, and uses the parameter-efficient fine-tuning method LoRA (Low-Rank Adaptation) for model optimization. By selecting 500 high-quality samples (the input is the original sample, and the output is the poisoned sample) for targeted training, we finally obtain a customized model that can stably generate high-quality data sets in specific fields. to generate backdoor samples, we can effectively avoid the above problems. For example, the poisoned sample generated by ”Dual-Trigger” is: ”Were he to aspire for both, he’d find himself achieving neither in the end.” This not only ensures the logical coherence between sentences, but also retains the core semantics of the original content, making the generated data fluent and natural without losing its intrinsic meaning. In Section 5.1, we will use an objective evaluation mechanism to comprehensively analyze the quality of poisoned datasets generated by different backdoor methods.
@@ 3.2.4 Poisoning rate setting and backdoor training
In backdoor attacks, the poisoning rate refers to the proportion of poisoned samples (samples with embedded triggers) injected into the training dataset. Different poisoning rates will significantly affect the effect of backdoor attacks. Backdoor attacks with low poisoning rates may be more covert and have less impact on model performance, but the ASR is low. Increasing the poisoning rate will correspondingly increase the ASR, but may significantly harm the CACC. Therefore, it is crucial to comprehensively compare the backdoor attack performance under different poisoning rates. In this paper, for the AG’s News dataset, the optimal poisoning rates of the three models are determined to be 5%, 10%, and 10%, respectively; while for the SST-2 and OLID datasets, the optimal poisoning rates are set to 30% and 20%, respectively.
@@ 4.1.1 Evaluation Datasets
This paper conducts in-depth research on three text classification tasks: sentiment analysis, offensive language identification, and news topic classification. The datasets used are Stanford Sentiment Treebank (SST-2) [32], Offensive Language Identification Dataset (OLID) [33], and AG’s News [34]. Table 2 lists the detailed information of the three datasets.
@@ 4.1.2 Victim Model
In recent years, pre-trained large language models (PLMs), represented by GPT-4 [9], have made breakthrough developments. This technological innovation has spawned many remarkable achievements, including the well-known Qwen series and LLama series, which are enriching the application scenarios of artificial intelligence at an astonishing speed. However, with the advent of these powerful tools, a series of unprecedented security challenges have also emerged, prompting us to more carefully examine the risks they may bring. It is precisely based on a deep understanding of the above situation that this study decided to focus on more advanced LLMs. In the end, we determined that the three victim models were Qwen2-72B-Instruct [35], LLama3-8B-Instruct [10], and LLama3.2-3B-Instruct [10], as shown in Table 3. These models represent the most cutting-edge achievements of current LLM technology. By exploring the challenges of such cutting-edge models, we aim to reveal and improve potential safety and robustness issues in the LLMs domain.
@@ 4.1.3 Baseline Methods
We select three representative textual backdoor attack methods as baseline methods. (1) BadNet [11], originally a backdoor attack method for images, was later cleverly adapted by Kurita and his team to a form suitable for textual attacks [14]. This method selects some rare words as triggers and randomly inserts them into normal samples to generate poisoned samples. (2) InsertSent [13], which uses fixed sentences as triggers and randomly inserts them into normal samples to generate poisoned samples. (3) Syntactic [15], which uses syntactic structures as triggers for textual backdoor attacks and conducts extensive experiments to demonstrate that the syntax-based attack method can achieve comparable attack performance to insertion-based methods, but with higher stealth and stronger defense capabilities.